OpenStudy (mendicant_bias):

Solving an IVP with Laplace-Heaviside, don't need help yet, just writing it out.

OpenStudy (mendicant_bias):

Using the Laplace Transform, solve the given Initial Value Problem. \[y'+2y=f(t), \ \ \ y(0)=0\]where\[f(t) =\begin{cases}t, \ \ \ 0 \leq t<1\\0, \ \ \ t\ge1.\end{cases}\]

OpenStudy (mendicant_bias):

First, putting the function f(t) in terms of a linear combination of Heaviside functions:\[f(t)=t-t \ \mathcal{U}(t-1).\] Plugging in and taking the Laplace transform of both sides,\[\mathcal{L}\left\{y'\right\}+\mathcal{L}\left\{2y\right\}=\mathcal{L}\left\{t-t \ \mathcal{U}(t-1)\right\} = sY(s)-y(0)+Y(s)=\frac{1}{s^2}-\frac{e^{-s}}{s^2}\]

OpenStudy (mendicant_bias):

\[Y(s)[s+2]=\frac{1}{s^2}-\frac{e^{-s}}{s^2},\]\[Y(s)=\frac{1-e^{-s}}{s^2(s-1)}\]

OpenStudy (mendicant_bias):

Splitting these up into two separate fractions and then taking partial fractions of the first (and applying it to the second one)-, I get\[\frac{1}{s^2(s-1)}=\frac{A}{s}+\frac{B}{s^2}+\frac{C}{(s+1)};\]\[As(s+1)+B(s+1)+Cs^2=1.\]

OpenStudy (sidsiddhartha):

uh is'nt it\[Y(s)=\frac{ 1-e^{-s}}{ s^2(s+2) }\]??

OpenStudy (mendicant_bias):

Yup, yup, my bad, totally missed the typo! Thank you.

OpenStudy (sidsiddhartha):

yeah and the rest is ok

OpenStudy (mendicant_bias):

\[s^2(A+C)+s(2A+B)+2B=1;\]\[B=\frac{1}{2},A=\frac{1}{4}, C=-\frac{1}{4}\]

OpenStudy (sidsiddhartha):

i'm getting something else

OpenStudy (sidsiddhartha):

:o

OpenStudy (mendicant_bias):

Lul one sec, lemme double check. \[s^2(A+C)=s^2(0); \ A = -C.\]\[s(2A+B)=s(0); \ 2A+B=0.\]OH, right there.

OpenStudy (mendicant_bias):

\[B=-\frac{1}{2}A\]

OpenStudy (sidsiddhartha):

yup

OpenStudy (mendicant_bias):

\[2B=1; \ B=\frac{1}{2}\]

OpenStudy (mendicant_bias):

Alright, so A = -1/4th, and C = 1/4.

OpenStudy (sidsiddhartha):

yes !!

OpenStudy (mendicant_bias):

That in mind, \[\Bigg[-\bigg(\frac{1}{4}\bigg)\frac{1}{s}+\bigg(\frac{1}{2}\bigg)\frac{1}{s^2}+\bigg(\frac{1}{4}\bigg)\frac{1}{s+2}\Bigg]\]

OpenStudy (mendicant_bias):

Alright, from there,\[-\frac{1}{4}\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s}\right\}+\frac{1}{2}\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s^2}\right\}+\frac{1}{4}\mathcal{L^{-1}}\left\{\vphantom{}\frac{1}{s+2}\right\}\]

OpenStudy (sidsiddhartha):

yes!!

OpenStudy (mendicant_bias):

(I feel like a huge troll taking my time to write all of the proper math, lol, sorry)

OpenStudy (usukidoll):

nah Latex is the right way to write it out

OpenStudy (sidsiddhartha):

\[\frac{ -1 }{ 4 }u(t)+\frac{ 1 }{ 2 }t.u(t)+\frac{ 1 }{ 4 }e^{-2t}u(t)\] seems fine?

OpenStudy (mendicant_bias):

\[-\frac{1}{4}+\frac{t}{2}+\frac{e^{-2t}}{4}\]Yeah, right now I'm just doing the first one but we can verbatim use the same inverse laplace for the second one with the unit step, so overall we should get

OpenStudy (sidsiddhartha):

yes but u can add u(t) with each term

ganeshie8 (ganeshie8):

Looks good to me \[f(t) = -\frac{1}{4} +\frac{t}{2}+\frac{e^{-2t}}{4}\] \[y(t) = f(t) - \mathcal{U}(t-1)f(t-1)\]

OpenStudy (mendicant_bias):

Oh my GOOOOODDD I just lost all that LaTeX written up due to a browser issue, one sec

OpenStudy (sidsiddhartha):

\[\frac{ e^{-s} }{ s^2(s+2) }=\frac{ -e^{-s} }{ 4 s }+\frac{ e^{-s} }{ 2s^2}+\frac{ e^{-s} }{ 4(s+2) }\]

OpenStudy (sidsiddhartha):

that ok?

OpenStudy (mendicant_bias):

Yuh

OpenStudy (sidsiddhartha):

yay

OpenStudy (sidsiddhartha):

u're a pro now u can do it :)

OpenStudy (mendicant_bias):

\[y(t)=-\frac{1}{4}+\frac{t}{2}+\frac{1}{4}e^{-2t}+\Bigg[-\frac{1}{4}+\frac{(t-1)}{2}+\frac{1}{4}e^{-(t-1)}\Bigg]\mathcal{U}(t-1)\]

OpenStudy (mendicant_bias):

Whoops, minus sign in between the first collection of terms and the second big term with the unit step

OpenStudy (sidsiddhartha):

yeah looks fine

OpenStudy (mendicant_bias):

And the exponent near the end, \[e^{-2(t-1)}\]

OpenStudy (mendicant_bias):

Alright, thank you!

OpenStudy (sidsiddhartha):

yeah !! :)

OpenStudy (sidsiddhartha):

really enjoy solving problems with you lulz :)

OpenStudy (mendicant_bias):

Lol, ty! You're very helpful.

OpenStudy (sidsiddhartha):

well i'm happy to be helpful :)